# Privacy-Preserving in Connected and Autonomous Vehicles Through Vision to Text Transformation

### v2

@@ 3 PROPOSED SOLUTION
By refining the output step by step, the system keeps only the essential information, removing redundant or irrelevant content. This not only reduces the volume of data transmitted to data centers but also ensures that each message precisely captures the corresponding visual scene. The formal definition of the RL model, incorporating its feedback-driven reward component, is outlined below.
Action Space: The action space comprises the selection of prompts from a predefined list at each stage. Each action corresponds to an index in the prompt list. For example, given 10 prompts in List A, the action space is a discrete set $\{0,1,\ldots,9\}$, where each index represents a specific prompt.
Observation Space: The observation space includes the current state of the generated text and any possible features extracted from it, such as embeddings generated by a language model.
Reward Set: The first reward is based on the relevance of the selected prompts to the currently generated text.
Feedback-Reward Set: The second reward is a feedback-based reward derived from RAG. As RAG often performs better assessments than the RL model’s internal reward signal, its output is used to perform an additional policy update.
Transition Set: The transition set consists of tuples $(s,a,r,s^{\prime})$ representing the interactions between states, actions, rewards, and resulting states.
@@ 4 RESULTS AND DISCUSSIONS
The experimental framework is implemented in Python 3.11, where input images are processed to generate descriptive text outputs. To evaluate the generalizability of the proposed model, two datasets are employed: Drive&Act [20], which contains approximately 9.6 million images, and a public traffic-scene dataset from Kaggle [21]. These datasets cover diverse environments to assess the model’s performance under varying conditions. Initially the model will be compared with the results obtained from a PPO-based RL model then it will be compared to the model introduced in [22]. Therefore, we also consider CFP-FP and AgeDB-30 datasets to apply .
The evaluation process involves two main categories of metrics: privacy preservation and text quality.
Table 2: Privacy Metrics Evaluation on the CFP-FP Dataset
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\downarrow$ SRRA (%) $\downarrow$
AdvFace 0.89 23.54 314.7 13.01
PPO-based RL 0.86 23.02 333.7 12.5
Feedback-based (Ours) 0.85 22.18 365.2 11.25
Table 3: Privacy Metrics Evaluation on the AgeDB-30 Dataset
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\downarrow$ SRRA (%) $\downarrow$
AdvFace 0.87 22.91 347.8 14.12
PPO-based 0.86 21.98 355.5 13.03
Feedback-based (Ours) 0.83 21.88 389.1 12.63
Figure 3: NER and Modifiers Metrics.
@@ 4.1 Privacy Preservation Metrics
To check how well privacy is protected, we use a test called the Adversarial Reconstruction Test. In this test, pictures are reconstructed and then compared against the original pictures. The following metrics are used where Table 2 and 3 presents results for the first four metrics, metrics in 5 are shown in Figure 3:
@@
    Named Entities and Modifiers evaluates whether the model correctly captures specific entities (e.g., people, places, objects) and uses rich descriptive modifiers. As illustrated in Figure  3, the baseline model often leaves such elements, while the proposed model consistently considers more.
@@ 4.2 Text Quality and Semantic Richness Metrics
The information and linguistic quality of the generated text is assessed by using the following literal and meaning metrics:
  • Word Count and Unique Word Count: reflect the length and literal variety of the output. As shown in Figure  4, the proposed model produces a longer and more detailed description than the baseline, and quantitatively indicates better expression.
  • Detailed Density: determines the amount of specific and descriptive information relative to the overall length of the text. It captures semantic prosperity by identifying constructions such as subordinate sections and detailed comments. Figure  5 highlights the proposed model produces intensive and more informative details, contributing to more accurate visual understanding.
@@
Figure 4: Lexical Richness metrics count the number of words.
Figure 5: Detailed Density is the amount of information per length.
@@
5 CONCLUSION AND FUTURE WORK
This study introduces a privacy-preserving framework for CAVs leveraging VLMs and RL. The proposed framework transforms visual data into concise, semantically rich text with the aim of preserving the individual privacy in the captured images. It demonstrates improved scalability, accuracy, and privacy protection, making it suitable for deployment in smart city infrastructures and industrial automation environments. Evaluations done both on privacy-related and text quality metrics confirm the effectiveness of the proposed method. Although the generated text captures key elements of the image, it excludes sensitive visual details, thereby preserving user privacy. The method can produce three to four times more descriptive content compared to baseline models, without compromising privacy. For future work, we propose exploring dynamic prompt generation, where the prompt list is adaptively updated at each iteration based on context or feedback. This would enable more flexible, personalized, and privacy-aware text generation, further enhancing the model’s generalizability and effectiveness.
ACKNOWLEDGMENTS
The work of Mehdi Sookhak was supported by the US Department of Transportation (USDOT), under Grant No. 69A3552348332, Tier-1 University Transportation Center (UTC) Transportation Cybersecurity Center for Advanced Research and Education (CYBER-CARE).

### v3

@@ 4 Proposed Solution
By refining the output step by step, the system keeps only the essential information, removing redundant or irrelevant content. This not only reduces the volume of data transmitted to data centers but also ensures that each message precisely captures the corresponding visual scene.
Reinforcement Learning Framework Details
To ensure clarity of the technical design, we elaborate the RL framework which coveres the state and action space formulations, reward function, exploration strategy, and training configuration. The RL agent operates in conjunction with the VLM and RAG module, guiding the text refinement process toward privacy-aware and semantically precise descriptions.
a) State Space Definition
The state space $S$ represents the environment information perceived by the RL agent at each iteration $t$. Each state $s_{t}$ encodes the semantic and privacy characteristics of the current textual output generated by the VLM:
$s_{t}=[E_{VLM}(x_{t}),\,P_{score}(x_{t})]$
where $E_{VLM}(x_{t})$ denotes the embedding vector obtained using the SBERT model for the generated text $x_{t}$, and $P_{score}(x_{t})$ is a privacy indicator derived from the dissimilarity between the original and reconstructed visual content (i.e., the inverse of visual similarity). This feature representation enables the agent to evaluate the text both semantically and in terms of privacy leakage potential.
b) Action Space Definition
The action space $A$ is defined as a discrete set of prompt-selection indices:
$A=\{a_{1},a_{2},...,a_{N}\}$
Each action $a_{t}\in A$ corresponds to selecting one textual prompt from the refined prompt list that guides the VLM to generate a new descriptive variant. Once the action is taken, the VLM produces a refined description $x_{t+1}$, and the environment transitions to a new state $s_{t+1}$ accordingly.
c) Reward Function Design
A composite reward function $R_{t}$ is designed to jointly optimize semantic privacy protection. The primary reward combines semantic similarity (measured by SBERT) and visual privacy (measured by reconstruction similarity):
$R_{t}=\alpha\cdot Sim_{SBERT}(x_{t},x_{gt})-\beta\cdot Sim_{VAE}(I_{orig},I_{rec})$
where $Sim_{SBERT}$ denotes the cosine similarity between the generated text $x_{t}$ and ground-truth or contextual text $x_{gt}$, and $Sim_{VAE}$ denotes the similarity between the original image $I_{orig}$ and its reconstructed version $I_{rec}$, estimated via a variational autoencoder (VAE). The coefficients $\alpha$ and $\beta$ control the trade-off between relevance and privacy.
To incorporate knowledge-based feedback, the reward is augmented with an external feedback term from the RAG module:
$R^{\prime}_{t}=R_{t}+\lambda\cdot F(s_{t},a_{t})$
where $F(s_{t},a_{t})$ represents the contextual validation score computed by RAG which indicates how strongly the retrieved information supports the generated text. The weighting factor $\lambda$ determines the influence of this feedback on the policy update.
d) Exploration–Exploitation Strategy
Proximal Policy Optimization (PPO) is employed to ensure stable policy learning while maintaining a balance between exploration and exploitation. The clipped surrogate objective function is given by:
$L_{PPO}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})}$ is the probability ratio between new and old policies, and $A_{t}$ is the advantage function defined as $A_{t}=Q(s_{t},a_{t})-V(s_{t})$. A decaying entropy term is added to encourage sufficient exploration during early training:
$L_{total}(\theta)=L_{PPO}(\theta)+\eta\cdot H[\pi_{\theta}(a|s)]$
where $H[\pi_{\theta}(a|s)]$ denotes the policy entropy and $\eta$ is gradually reduced (from $0.01$ to $0.001$) every 50 epochs. The clipping parameter $\epsilon$ is set to $0.2$ to maintain stable updates and prevent overshooting.
e) Network Architecture and Hyperparameters
The policy and value networks are implemented using two fully connected layers of sizes $512$ and $256$, followed by ReLU activations. The output layer corresponds to the size of the action space $|A|$. The optimizer is Adam with a learning rate of $3\times 10^{-5}$, and training proceeds for $500$ iterations using a batch size of $32$. Convergence is determined when both the cumulative reward and the semantic similarity score stabilize. The training pipeline is summarized as follows:
  • 1.
    Initialize VLM parameters and prompt list.
  • 2.
    Encode initial text and compute:
    $s_{t}=[E_{VLM}(x_{t}),P_{score}(x_{t})]$.
  • 3.
    Select action $a_{t}$ based on policy $\pi_{\theta}(a|s_{t})$.
  • 4.
    Generate refined text $x_{t+1}$ and compute $R_{t}$, $R^{\prime}_{t}$.
  • 5.
    Update policy using $L_{total}(\theta)$.
This configuration ensures stable convergence while progressively improving both privacy protection and semantic richness of generated descriptions.
@@ 5 Results and Discussions
The experimental framework is implemented in Python 3.11, where input images are processed to generate descriptive text outputs. To evaluate the generalizability of the proposed model, two datasets are employed: CFP-FP [14] dataset which is a face verification benchmark containing celebrity images in both frontal and profile views, designed to evaluate pose-invariant face recognition performance. Additionally, AgeDB-30 which is used for face verification under large age gaps, typically involving 570 identities and  12,240 images. The model will be compared with the results obtained from a VLM-based, RL-based, and feedback-based RL model then it will be compared to the model introduced in [16] and [11].
The evaluation process involves four main categories of evaluations: Privacy Preservation Metrics, Privacy Quantification and Risk Analysis, Text Quality and Semantic Richness Metrics, Ablation Study.
A sample textual description generated by the proposed model is shown in Figure 5 which is compared with the output from two other models. Then the textual description generated from the proposed model is used to generate a new image and it is shown to be significantly different from the input image.
Figure 5: The figure illustrates sample textual descriptions generated by the proposed model and two additional models. The first description is produced by a VLM-based model, while the second is generated by a VLM-RL model. The third description is produced by the proposed model, which is capable of generating significantly longer and more detailed textual outputs. Additionally, the red-colored highlights indicate textual richness, showing that the proposed model effectively captures the main elements of the image. Using the generated text, an text to image model was used to evaluate how close is the reconstructed image to the input image in terms of information leakage. It is obvious that the re-generated image is a vehicle but without any similarity to the input image, meanwhile safety.
@@ 5.1 Privacy Preservation Metrics
To check how well privacy is protected, we use a test called the Adversarial Reconstruction Test. In this test, pictures are reconstructed and then compared against the original pictures. The following metrics are used where Table 3 and 4 presents results for the first four metrics, metrics in 5 are shown in Figure 6:
@@
    Named Entities and Modifiers evaluates whether the model correctly captures specific entities (e.g., people, places, objects) and uses rich descriptive modifiers. As illustrated in Figure  6, the baseline model often leaves such elements, while the proposed model consistently considers more.
Table 3: Privacy Metrics Evaluation on the CFP-FP Dataset
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\uparrow$ SRRA (%) $\downarrow$
AdvFace [16] 0.89 23.54 314.7 13.01
FLDATN [11] 0.91 24.32 349.1 -
PPO-based RL 0.86 23.02 333.7 12.5
Feedback-based RL (Ours) 0.85 22.18 365.2 11.25
Table 4: Privacy Metrics Evaluation on the AgeDB-30 Dataset
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\uparrow$ SRRA (%) $\downarrow$
AdvFace [16] 0.87 22.91 347.8 14.12
FLDATN [11] 0.89 24.59 369.1 -
PPO-based 0.86 21.98 355.5 13.03
Feedback-based RL (Ours) 0.83 21.88 389.1 12.63
Table 5: Differential Privacy Trade-Off Between Privacy Budget and Text Quality
Noise Level $\sigma$ $\varepsilon$ (↓) BLEU Score (↓) Privacy Gain (% ↑)
0.0 $\infty$ 0.831 0
0.1 5.3 0.782 9.1
0.2 3.1 0.779 15.7
0.3 2.2 0.756 23.4
5.2 Privacy Quantification and Risk Analysis
To establish a rigorous and quantifiable evaluation of privacy preservation, the proposed Hierarchical Feedback RL-VLM framework is complemented by formal privacy metrics and empirical risk analysis. Differential Privacy (DP) evaluation is performed to measure the privacy leakage and verify that textual transformation suppresses identity-linked cues rather than just removing visual content.
Differential Privacy Analysis
In order to provide a formal privacy guarantee, a DP mechanism is applied to the sentence-embedding layer of the VLM. Gaussian noise $\mathcal{N}(0,\sigma^{2})$ is injected into each textual embedding vector before storage or transmission:
$\tilde{E}_{VLM}(x_{t})=E_{VLM}(x_{t})+\mathcal{N}(0,\sigma^{2})$
According to the Gaussian mechanism, the perturbed embedding satisfies $(\varepsilon,\delta)$-DP, where
$\varepsilon=\frac{\Delta_{2}f}{\sigma},\quad\delta<10^{-5}$
and $\Delta_{2}f$ is the $\ell_{2}$-sensitivity of the embedding function. By varying the noise scale $\sigma$, the privacy utility trade-off can be measured. A higher $\sigma$ ensures stronger privacy (smaller $\varepsilon$) at the cost of reduced linguistic fidelity.
Table 5 illustrates that adding small Gaussian noise significantly reduces the privacy budget $\varepsilon$ with minimal degradation in text quality which establishes a measurable privacy utility balance.
@@ 5.3 Text Quality and Semantic Richness Metrics
The information and linguistic quality of the generated text is assessed by using the following literal and meaning metrics to evaluate the quality of privacy-preserved generated textual description:
  • 1.
    Word Count and Unique Word Count: reflect the length and literal variety of the output. As shown in Figure  8, the proposed model produces a longer and more detailed description than the baseline, and quantitatively indicates better expression.
  • 2.
@@
Figure 6: NER and Modifiers Metrics.
@@
Word CountUnique Word Count0100100200200300300747446461311318989223223149149CountVLMRL-VLMProposed Method
Figure 8: Lexical Richness metrics count the number of words such the proposed method outperforms other models.
5.4 Ablation and Comparative Study
To understand the contribution of each component in the proposed hierarchical feedback-based RL-VLM model, an ablation study was conducted. Four model variants were evaluated by enabling (i) hierarchical RL and (ii) feedback-based retrieval using RAG. All models were tested on the same experimental setup described earlier.
Table 6: Model Configurations Used in the Ablation Study
Model Variant VLM Hierarchical RL Feedback (RAG)
Baseline (VLM Only) $\times$ $\times$
RL-VLM (Single - Stage) $\times$
RL-VLM + RAG (Feedback)
Full Model (Hierarchical Feedback RL-VLM) ✓✓ ✓✓ ✓✓
Figure 9 presents the quantitative evaluation across both privacy and text-quality metrics. Performance consistently improves as each component is integrated, demonstrating the significance of hierarchical learning and feedback refinement.
Qualitative Comparison
Figure 9 illustrates the qualitative differences between the four variants for the same visual input. The VLM-only baseline produces generic one-line captions with limited privacy awareness, while the RL-VLM variant adds contextual structure but lacks external correction. Integrating RAG feedback leads to semantically richer text and the full Hierarchical Feedback RL-VLM obtains the most descriptive and privacy-aware output.
Discussion
The ablation results demonstrate that:
  1. 1.
    Hierarchical RL substantially enhances semantic precision and fine-grained scene understanding by guiding prompt refinement across multiple iterations.
  2. 2.
    Feedback via RAG introduces external semantic correction, improving linguistic diversity and alignment with privacy objectives.
  3. 3.
    The combined iterative framework achieves the best overall performance, significantly outperforming traditional single-pass VLM captioning and single-stage RL baselines.
Overall, the experiments validate that the proposed hierarchical feedback-based reinforcement learning framework extends beyond prior captioning and RL-VLM approaches, providing a systematic and adaptive mechanism for privacy-preserving visual-to-text transformation.
SSIM ($\downarrow$)MSE ($\uparrow$)SRRA ($\downarrow$)Unique Words ($\uparrow$)Detail Density ($\uparrow$)Semantic Score ($\uparrow$)00.50.5110.10.10.150.150.350.350.220.220.40.40.330.330.30.30.230.230.430.430.280.280.550.550.620.62111111111111Normalized Score (0–1)VLM OnlyRL-VLMFull (HFR-VLM)
Figure 9: Ablation Study Comparison of Model Variants.
Conclusion and Future Work
This study introduces a privacy-preserving framework for ITS leveraging VLMs and RL. The proposed framework transforms visual data captured in ITS into concise, semantically rich text with the aim of preserving the individual privacy in the captured images. It demonstrates improved scalability, accuracy, and privacy protection, making it suitable for deployment in smart city infrastructures and industrial automation environments. Evaluations performed both on privacy-related and text quality metrics confirm the effectiveness of the proposed method. Although the generated text captures key elements of the image, it excludes sensitive visual details, thereby preserving user privacy. The method can produce three to four times more descriptive content compared to baseline models, without compromising privacy. For future work, we propose exploring dynamic prompt generation, where the prompt list is adaptively updated at each iteration based on context or feedback. This would enable more flexible, personalized, and privacy-aware text generation, further enhancing the model’s generalizability and effectiveness.